### v1

# Illuminating Scalar Dark Matter Co-Scattering in EFT with Monophoton Signatures

Abstract
We investigate the co-scattering mechanism for dark matter production in an EFT framework which contains new $Z_{2}$-odd singlets, namely two fermions $N_{1,2}$ and a real scalar $\chi$. The singlet scalar $\chi$ is the dark matter candidate. The dimension-5 operators play a vital role to set the observed DM relic density. We focus on a nearly degenerate mass spectrum for the $Z_{2}$ odd particles to allow for a significant contribution from the co-scattering or co-annihilation mechanisms. We present two benchmark points where either of the two mechanisms primarily set the DM relic abundance. The main constraint on the model at the LHC arise from the ATLAS mono-$\gamma$ search. We obtain the parameter space allowed by the observed relic density and the mono-$\gamma$ search after performing a scan over the key parameters, the masses $M_{N_{1,2}},M_{\chi}$ and couplings $c_{3}^{\prime},Y^{\prime}_{11,22}$. We find the region of parameter space where the relic abundance is set primarily by the co-scattering mechanism while being allowed by the LHC search. We also determine how the model can be further probed at the HL-LHC via the mono-$\gamma$ signature.
Keywords:
Beyond the Standard Model, Particle Nature of Dark Matter, Dark Matter at Colliders
preprint: CQUeST-2025-0756
IP/BBSR/2025-02
1 Introduction
Despite strong evidence for dark matter in astrophysics and cosmology, the nature of dark matter (DM) remains unknown Bertone:2004pz ; Cirelli:2024ssz . Moreover searches for new physics at the high energy or intensity frontiers have not provided clear evidence for physics beyond the standard model. This leaves a large number of possibilities for what dark matter might be. Traditionally the new physics models that were proposed to provide a dark matter candidate and address open issues with the standard model (SM), with neutrino mass, or with baryon asymmetry propose a new weakly interacting particle that is in thermal equilibrium with the standard model bath in the early Universe and freezes-out as the Universe cools down Arcadi:2017kky . The absence of signals in various astroparticle and collider searches for WIMPs has rekindled the interest in alternative mechanisms for DM formation and include various modifications of the freeze-out process such as secluded freeze-out Pospelov:2007mp ; Feng:2008mu , DM conversion or co-scattering Garny:2017rxs ; DAgnolo:2017dbv , forbidden annihilation DAgnolo:2015ujb , 3-2 annihilation Hochberg:2014dra , assisted freeze-out Belanger:2011ww and interactions between dark sectors Belanger:2012vp ; Belanger:2021lwd , decoupled freeze-out Chu:2011be or processes where thermal equilibrium is not achieved such as in the freeze-in mechanism McDonald:2001vt ; Hall:2009bx or with axion dark matter Adams:2022pbo . These expand the range of possibilities for dark matter masses and interaction strengths as well as for the content of the dark sector thus impacting the particle and astroparticle physics signatures that can be expected. For example when dark matter is feebly interacting and freezes-in collider signatures are linked to the presence of long-lived particles Feng:2003uy ; Belanger:2018sti ; Belanger:2022gqc .
In the DM conversion or coscattering mechanism, DM is weakly interacting but the WIMP self-annihilation is suppressed such that the process that drives DM formation relies on the interaction between DM and another almost degenerate particle in the dark sector Garny:2017rxs ; DAgnolo:2017dbv . Either coannihilation of a DM particle with another particle of the dark sector Binetruy:1983jf ; Griest:1990kh or co-scattering of the DM with the SM into another particle of the dark sector can be the dominant process. These processes typically involve weaker couplings of DM to the SM than in the freeze-out scenario although couplings are assumed to be large enough for DM to be in equilibrium with the SM bath. Several models that rely on this mechanism were studied  Cheng:2018vaj ; Brummer:2019inq ; Junius:2019dci ; Garny:2021qsr ; Alguero:2022inz ; Heeck:2022rep ; Heisig:2024mwr ; Heisig:2024xbh ; DiazSaez:2024dzx ; DiazSaez:2024nrq and some signatures involving long-lived particles were explored Alguero:2022inz ; DiazSaez:2024dzx . In this paper we study the co-scattering mechanism in a simplified extension of the SM which contains two singlet fermions and a new real singlet scalar which we assume to be the dark matter. We work within an EFT framework and include interactions through dimension 5 operators. Such model was considered in  Belanger:2021slj in the regime where DM is feebly interacting and its interactions are linked to the generation of neutrino masses.
In this work we explore the regions of parameter space where the singlet and lightest fermion are nearly degenerate in mass such that either co-scattering or coannihilation processes dominate dark matter formation. We assume that the quartic coupling between the scalar singlet and the Higgs is small to evade the strong bounds from direct detection  Yu:2024xsy ; GAMBIT:2017gge , hence the quartic interactions do not play a role in DM formation. Co-scattering is typically dominated by the decay of the singlet fermion to DM and a neutrino. The DM relic density is obtained after solving two Boltzmann equations for the fermion and the scalar respectively. Because of the small couplings involved, the main signature is not directly linked with DM but results from the production and decay of the heavy fermion into the lightest fermion and a photon. The signature is challenging because of the low energy of the photon when the two fermions have a small mass splitting. We revisit current LHC constraints arising from the ATLAS monophoton search ATLAS:2020uiq and make projections for the HL-LHC. We show that some parameter space of the model can be probed at the LHC.
The structure of this paper is as follows: The model is described in Section 2 , Section 3 discusses DM formation and describes the computation of the relic density. Section 4 presents the analysis relevant for establishing the LHC constraints. Section 5 is devoted to numerical results including both DM and collider observables. Finally, we conclude with a summary in Section 6.
2 The Model
We propose an EFT framework where the SM field contents is extended with two SM singlet fermions $N_{1},N_{2}$ and a real singlet scalar $\chi$. We impose a $Z_{2}$ symmetry under which both $N_{i}$ and $\chi$ are odd, and all the SM fields are even. We consider $\chi$ as the DM candidate.
The Lagrangian of this EFT framework up-to d=5 operators is written as
$\mathcal{L}_{eff}$ $=$ $\mathcal{L}_{SM}+\frac{c_{5}}{\Lambda}(\overline{L^{c}}\tilde{\Phi})(\tilde{\Phi}^{\dagger}L)+M_{B}\overline{N^{c}}N+\frac{Y}{\Lambda}\overline{L}\tilde{\Phi}N\chi$
$+\frac{c_{1}}{\Lambda}\overline{N^{c}}N\chi^{2}+\frac{c_{2}}{\Lambda}\overline{N^{c}}N\Phi^{\dagger}\Phi+\frac{c_{3}}{\Lambda}\overline{N^{c}}\sigma_{\mu\nu}NB^{\mu\nu}+h.c.$
$+\lambda\Phi^{\dagger}\Phi\chi^{2}+\beta\chi^{4}.$
In the above $\tilde{\Phi}=i\sigma_{2}\Phi^{\star}$ and $N^{c}=(i\gamma^{2}\gamma^{0})\overline{N}^{T}$, where $N=(N_{1},N_{2})^{T}$. Note that in the above Lagrangian flavor indices are implicit. The couplings $c_{5}$ is a $3\times 3$ matrix, $Y$ is a $3\times 2$ matrix and $c_{1,2,3}$ are $2\times 2$ matrices in flavor space. The dipole operator, $\frac{c_{3}}{\Lambda}N^{T}C^{-1}\sigma_{\mu\nu}NB^{\mu\nu}$, plays an important role in the phenomenology, in particular it leads to the decay of $N_{2}\to N_{1}\gamma$ which will give the dominant signature of the model at the LHC. This operator is relevant only when considering at least two singlet fermions, which is why we introduced two singlet fermions. Note that the renormalizable Yukawa interaction $\overline{L}\Phi N$ is forbidden due to the $Z_{2}$ symmetry. Therefore, the field $N_{i}$ can not participate in the SM neutrino mass generation at tree level via the d=5 Weinberg operator and there is no mixing between SM neutrino and $N_{i}$. After EW symmetry breaking, $c_{2}$ contribute to the mass of $N$ as $M_{N}=M_{B}+c_{2}v^{2}/\Lambda$. We redefine the couplings of the d=5 operators as, $c_{5}^{\prime}=\frac{c_{5}}{\Lambda},Y^{\prime}=\frac{Y}{\Lambda},c_{1}^{\prime}=\frac{c_{1}}{\Lambda},c_{2}^{\prime}=\frac{c_{2}}{\Lambda},c_{3}^{\prime}=\frac{c_{3}}{\Lambda}$. For simplicity, we consider all the Wilson co-efficient $c_{i}^{\prime}$ and $Y^{\prime}$ to be diagonal except for $c_{3}^{\prime}$ to be off-diagonal.
We have explored the FIMP dark matter in a similar model which however featured a different discrete symmetry  Belanger:2021slj . In that model the additional $\mathcal{Z}_{2}$ symmetry (that ensures stability of the DM particle $N_{3}$) forbids the d=5 dipole operator.
Figure 1: Schematic diagrams representing different WIMP $\chi$ dilution mechanisms in the early Universe for our model.
3 Dark Matter Freeze-Out
We consider both $\chi$ and $N_{1,2}$ to be in thermal equilibrium with the SM plasma. The thermal equilibrium of $\chi$ is maintained by the following processes,
  • $\chi$ pair annihilation: $\chi\chi\to\text{SM}\leavevmode\nobreak\ \text{SM}$.
  • $\chi$ and $N_{1,2}$ co-annihilation: $\chi\leavevmode\nobreak\ N_{1,2}\to\text{SM}\leavevmode\nobreak\ \text{SM}$.
  • $N_{1,2}/\chi$ interchange: $N_{1,2}\leavevmode\nobreak\ \text{SM}\to\chi\leavevmode\nobreak\ \text{SM}$, $\chi\leavevmode\nobreak\ \nu_{e,\mu}\to N_{1,2}$ .
Furthermore, the thermal equilibrium of $N_{1,2}$ is maintained either by the same interactions or by the following processes,
  • $N_{1,2}$ pair annihilation: $N_{1,2}\leavevmode\nobreak\ N_{1,2}\to\gamma\gamma,\gamma Z,ZZ$.
  • $N_{1}$ and $N_{2}$ co-annihilation: $N_{1}\leavevmode\nobreak\ N_{2}\to f\bar{f},W^{+}W^{-},ZH$.
We assume $\chi$ to be the DM particle and $N_{1,2}$ the next-to-lightest-odd (NLOP) particles. We demand $M_{{N}_{1,2}}<1.5M_{\chi}$ to ensure a significant contribution from either co-annihilation Klasen:2013jpa ; Molinaro:2014lfa or co-scattering DAgnolo:2017dbv ; Junius:2019dci processes with $\chi$ states. Hence, the abundance of DM $\chi$ is governed by the interplay of $N_{1,2}$ and $\chi$ interactions in the early Universe. Schematic diagrams representing different $\chi$ dilution in the early Universe are shown in Fig. 1, in addition $\chi\nu_{e,\mu}\to N_{1,2}$ contribute to co-scattering.
For convenience and to keep with the formalism used in micrOMEGAs Alguero:2022inz we separate all particles into three sectors, namely, sector 0 corresponding to SM particles, sector $1$ comprising $N_{1}$ and $N_{2}$, and sector $2$ containing $\chi$. The particles within each sector are assumed to be in thermal equilibrium. The evolution of the dark sector particles in the early Universe is governed by the Boltzmann equations Alguero:2022inz ,
$\frac{dY_{1}}{dx}$ $=$ $-\frac{1}{x^{2}}\frac{s(M_{\chi})}{\tilde{H}(M_{\chi})}\left[\langle\sigma_{1100}v\rangle(Y_{1}^{2}-{Y_{1}^{eq}}^{2})+\langle\sigma_{1122}v\rangle\left(Y_{1}^{2}-Y_{2}^{2}\frac{{Y_{1}^{eq}}^{2}}{{Y_{2}^{eq}}^{2}}\right)\right$. (2)
$+\left.\langle\sigma_{1200}v\rangle(Y_{1}Y_{2}-Y_{1}^{eq}Y_{2}^{eq})-\frac{\Gamma_{2\rightarrow 1}}{s}\left(Y_{2}-Y_{1}\frac{Y_{2}^{eq}}{Y_{1}^{eq}}\right)\right]\,$,
$\frac{dY_{2}}{dx}$ $=$ $-\frac{1}{x^{2}}\frac{s(M_{\chi})}{\tilde{H}(M_{\chi})}\left[\langle\sigma_{2200}v\rangle(Y_{2}^{2}-{Y_{2}^{eq}}^{2})-\langle\sigma_{1122}v\rangle\left(Y_{1}^{2}-Y_{2}^{2}\frac{{Y_{1}^{eq}}^{2}}{{Y_{2}^{eq}}^{2}}\right)\right$. (3)
$+\left.\langle\sigma_{1200}v\rangle(Y_{1}Y_{2}-Y_{1}^{eq}Y_{2}^{eq})+\frac{\Gamma_{2\rightarrow 1}}{s}\left(Y_{2}-Y_{1}\frac{Y_{2}^{eq}}{Y_{1}^{eq}}\right)\right]\,.$
Here $Y_{1,2}$ is the abundance of the particles in the dark sector 1 and 2, and $Y_{1,2}^{eq}$ is the corresponding equilibrium abundance. $\langle\sigma_{ijkl}v\rangle$ indicates the thermal average cross-section for the process $x_{i}x_{j}\to x_{k}x_{l}$ where $x_{i}$ stands for any particle in sector i. The term $\Gamma_{2\to 1}$ is the conversion term, including both co-scattering as well as decay terms, and is given by,
$\Gamma_{2\rightarrow 1}=\frac{\sum_{\alpha=1}^{2}\Gamma_{N_{\alpha}\to\chi,\nu_{e,\mu}}\,g_{\alpha}M_{N_{\alpha}}^{2}K_{1}\left(\frac{M_{N_{\alpha}}}{M_{\chi}}x\right)}{\sum_{\alpha=1}^{2}g_{\alpha}M_{N_{\alpha}}^{2}K_{2}\left(\frac{M_{N_{\alpha}}}{M_{\chi}}x\right)}+\langle\sigma_{2010}v\rangle s(M_{\chi}/x)Y^{eq}_{0}\,$, (4)
where $x=M_{\chi}/T$, $g_{\alpha}(=2)$ is the internal degree of freedom of $N_{1,2}$, $K_{1},K_{2}$ are the modified Bessel function of the second kind of order one and two and $Y^{eq}_{0}(=0.238)$ is the SM sector comoving equilibrium number density. Note that the modified Hubble rate and entropy density in terms of $M_{\chi}$ is given by,
$\tilde{H}(M_{\chi})=H(M_{\chi})\left[1-\frac{x}{3g^{s}_{*}}\frac{dg^{s}_{*}}{dx}\right]^{-1},\,\,\,\,\,\,s(M_{\chi})=\frac{2\pi^{2}}{45}g^{s}_{*}M_{\chi}^{3},$ (5)
where $g^{s}_{*}$ is the effective number of degrees of freedom related to the entropy density of the Universe, $s$.
We used micrOMEGAs Alguero:2022inz ; Alguero:2023zol to solve Eq. 2 and Eq. 3. For this we implemented the model in Feynrules Alloul:2013bka and generated the model files for CalcHEP Belyaev:2012qa required by micrOMEGAs. The code then solves Eq. 2 and 3 for the abundances and the relic density given by,
$\Omega_{\chi}h^{2}=2.742\times 10^{8}(M_{N_{1}}Y_{1}+M_{\chi}Y_{2}).$ (6)
In our analysis, we assume $m_{N_{2}}>m_{N_{1}}$, and thus $N_{2}$ decouples prior to $N_{1}$ in sector 1. Consequently, at the late epoch of the Universe, only $N_{1}$ contributes to $Y_{1}$. The NLOP $N_{1}$ decays to DM $\chi$ after $N_{1}$ freeze-out, i.e $Y_{1}\to 0$ at $T<<M_{N_{1,2}}$, such that the only relevant term that contributes to the observed relic density is from the second term $Y_{2}$ in Eq. 6.
The DM relic density is precisely determined by the PLANCK satellite through measurements of the cosmic microwave background (CMB) Planck:2018vyg ,
$\Omega_{\chi}h^{2}=0.12\pm 0.0012.$ (7)
In the numerical analysis we will require that the DM relic density lies within $2\sigma$ of the observed value.
Figure 2: Feynman diagrams for the conversion processes that depend on the Yukawa coupling $Y^{\prime}$. The inverse decay shown in the bottom line usually dominates.
Initial state Final state Scaling with couplings
$\chi$ $h$ $N_{1,2}$ $\nu_{e,\mu}$ ${Y^{{}^{\prime}}}_{11(22)}^{2}$
$\chi$ $W^{\pm}$ $N_{1,2}$ $e^{\pm}\mu^{\pm}$ ${Y^{{}^{\prime}}}_{11(22)}^{2}$ (t-channel process )
$\chi$ $\nu_{e,\mu}$ $N_{1,2}$ - ${Y^{{}^{\prime}}}_{11(22)}^{2}$ (Inverse Decay)
Table 1: List of all relevant conversion processes shown in Fig. 2. All cross sections depend on the square of the Yukawa coupling $Y^{\prime}$.
To illustrate the importance of the co-scattering and co-annihilation mechanisms for obtaining the correct relic density, we introduce $\Delta^{1}_{\chi}$, which is the fractional difference between the relic densities obtained after solving the coupled and single Boltzmann equation Alguero:2022inz :
$\Delta^{1}_{\chi}\equiv 1-\frac{\Omega h^{2}(\textrm{Single})}{\Omega h^{2}(\textrm{Coupled})}.$ (8)
More specifically, $\Omega h^{2}(\text{Single})$ refers to the value obtained using the standard darkOmega function in micrOMEGAs, which accounts for only one Boltzmann equation and thus incorporates solely co-annihilation processes. In contrast, $\Omega h^{2}(\text{Coupled})$ represents the result obtained via the darkOmegaN function, where the dark particles are divided into two sectors. This calculation involves solving the coupled system of Boltzmann equations 2 and 3, including all co-scattering and decay processes.
Figure 3: Feynman diagrams for the pair annihilation and co-annihilation process of $N_{1,2}$ that depend on the coupling $c_{3}^{\prime}$. Note that due to the choice $c^{\prime ij}_{3}=\epsilon^{ij}c^{\prime ij}_{3}$, we have both co-annihilation and annihilation processes between $N_{1}$ and $N_{2}$.
Initial state Final state Scaling with couplings
$N_{1,2}$ $N_{1,2}$ $\gamma,Z$ $\gamma,Z$ ${c_{3}^{{}^{\prime}}}^{4}$ (t- channel process )
$N_{1}$ $N_{2}$ $f$ $\bar{f}$ ${c_{3}^{{}^{\prime}}}^{2}$ (s-channel process )
$N_{1}$ $N_{2}$ $^{+} WW^{-}$ ${c_{3}^{{}^{\prime}}}^{2}$ (s-channel process )
$N_{1}$ $N_{2}$ $Z$ $H$ ${c_{3}^{{}^{\prime}}}^{2}$ (s-channel process )
Table 2: List of all relevant annihilation and co-annihilation processes as in Fig. 3, with the dependency of their cross sections on the $c_{3}^{\prime}=c_{3}^{{}^{\prime}12}$.
Note that both sectors remain in thermal equilibrium when the Wilson coefficient ($c_{i}^{{}^{\prime}},\,Y^{{}^{\prime}}$) and the Higgs portal coupling $\lambda$ are sufficiently large. If this is the case $Y_{1}/Y_{2}=Y_{1}^{eq}/Y_{2}^{eq}$ and Eq. 23 simplify considerably and solving the coupled Boltzmann equations yields the same result as solving a single abundance equation where $Y_{1}=0$ and $Y_{2}=Y$. Hence, the usual freeze-out result is recovered and $\Delta^{1}_{\chi}=0$. Conversely, in scenarios where $Y^{{}^{\prime}},\lambda\sim\mathcal{O}(10^{-6}-10^{-10})$, $\chi$’s pair annihilation, co-annihilation, and exchange process involving $\chi$ and constituents of sector 2 become negligible, thereby simplifying the Boltzmann equations [^1] For the numerical analysis, we solve the general Boltzmann equations governing the evolution of $\chi$ and the constituents of sector 1, specifically Eqs. 2 and 3.,
$\frac{dY_{1}}{dx}=-\frac{1}{x^{2}}\frac{s(M_{\chi})}{\tilde{H}(M_{\chi})}\left[\langle\sigma_{1100}v\rangle(Y_{1}^{2}-{Y_{1}^{eq}}^{2})-\frac{\Gamma_{2\rightarrow 1}}{s}\left(Y_{2}-Y_{1}\frac{Y_{2}^{eq}}{Y_{1}^{eq}}\right)\right]\,$, (9)
$\frac{dY_{2}}{dx}=-\frac{1}{x^{2}}\frac{s(M_{\chi})}{\tilde{H}(M_{\chi})}\left[\frac{\Gamma_{2\rightarrow 1}}{s}\left(Y_{2}-Y_{1}\frac{Y_{2}^{eq}}{Y_{1}^{eq}}\right)\right]\,.$ (10)
The dominant process in the above equation contributing to $\Gamma_{2\to 1}$ arises from $N_{1,2}$ decay while $2\to 2$ processes are subdominant. The corresponding Feynman diagrams are presented in Fig. 2. All cross-sections depend on ${Y^{{}^{\prime}}}^{2}$ as shown in Table 1. It is important to emphasize that we need at least $N_{1}$ to be in chemical equilibrium with the thermal bath through (co-)annihilation process, i.e. 1100, which depends on $c_{3}^{\prime}$ solely, to facilitate $\chi$ dilution either through co-annihilation or co-scattering. Therefore, it becomes necessary to assume $c_{3}^{\prime}>10^{-5}$ to ensure that the constituents of sector 1 remain in chemical equilibrium at $T\geq M_{N_{1,2}}$. On the contrary, due to the small coupling of $\chi$ with $N_{1,2}$ in this scenario, $\chi$ may undergo chemical decoupling at much larger $T$. The relic density of $\chi$ is set when the conversion or co-scattering processes shut off and it freezes out. Furthermore, as we focus on the conversion/co-annihilation processes, it is convenient to define the following variables, which dictate the relative mass splitting between the dark sector particles,
$\delta_{1}=\frac{{M_{N_{1}}}-M_{\chi}}{M_{\chi}},\,\,\,\,\,\,\delta_{2}=\frac{{M_{N_{2}}}-{M_{N}}_{1}}{{M_{N_{1}}}}$ (11)
Before focusing on the primary analysis of this work, we want to bring to the readers’ attention that $\chi$ pair annihilation through the Higgs portal coupling $\lambda$ is stringently constrained from direct detection (DD) experiments such as LUX-Zeplin Lz:2024zvo . Detailed analysis of $\chi$ pair annihilation analysis through the Higgs portal coupling $\lambda$ can be found in  Escudero:2016gzx ; GAMBIT:2017gge ; Yu:2024xsy ; Belanger:2024wca ; Escudero:2025eid . To evade the stringent constraint on $\lambda$ from DD searches while ensuring that $\chi$ remains in thermal equilibrium, we fix $\lambda$ to $10^{-6}$ such that,
$\left(\frac{\Gamma_{\chi\chi\to h\to\text{SM}\text{SM}}}{H}\right)_{T=M_{\chi}}\simeq\mathcal{O}(1-100),$ (12)
where $\Gamma_{\chi\chi\to h\to\text{SM}\text{SM}}$ is the $\chi$ pair annihilation rate in the early Universe through the Higgs portal coupling $\lambda$.
In this work, we focus primarily on the thermal production of $\chi$ in a scenario where $c_{3}^{\prime}$ and $Y^{\prime}$ are non-zero, while fixing $\lambda=10^{-6}$ and setting all other Wilson coefficients to zero [^2] Assuming $c_{1,2}\neq 0$, $Y^{\prime}\neq 0$, and $c_{3}^{\prime}=0$, similar DM phenomenology may arise under the current assumptions of this work. A thorough exploration of additional scenarios is deferred to future work.. In this case, the value of $\delta_{2}$ determines the dominant process contributing to $N_{1}$ annihilation. When $\delta_{2}<0.5$, $N_{1}$ is primarily diluted through co-annihilation while pair annihilation is subdominant; otherwise, pair annihilation of $N_{1}$ is the dominant process Masso:2009mu ; Weiner:2012cb ; Tulin:2012uq ; Cline:2012bz . Note that the dominant co-annihilation process is $N_{1}N_{2}\to f\bar{f}$. In comparison, the process $N_{1}N_{2}\to W^{+}W^{-}$ contributes less than 1% to the total co-annihilation rate due to a cancellation between the virtual $Z$ and $\gamma$ exchange diagrams. The relevant Feynman diagrams for $N_{1}$ dilution in the early Universe are shown in Fig. 3, and the interaction rates dependence on $c_{3}^{{}^{\prime}}$ are given in Table 2.
We begin by examining two benchmark points to illustrate how the evolution of DM $\chi$ depends on the key parameters, i.e. $c_{3}^{\prime},Y^{\prime},m_{\chi}\,\text{and}\,m_{N_{1,2}}$. A detailed analysis of the allowed parameter space will be presented in Section 5. Based on the primary production mechanism of $\chi$, either through co-annihilation or co-scattering, we have considered two benchmark points. We assume $c_{3}^{\prime}=7\times 10^{-4}\leavevmode\nobreak\ \text{GeV}^{-1}$ and the other parameters are set to,
  • BP1: $M_{\chi}=100\,\text{GeV},\,M_{N_{1}}=105\,\text{GeV},\,M_{N_{2}}=115\,\text{GeV},\,{Y^{\prime}}_{11}=Y^{\prime}_{22}=1.1\times 10^{-9}\leavevmode\nobreak\ \text{GeV}^{-1}$
  • BP2: $M_{\chi}=300\,\text{GeV},\,M_{N_{1}}=315\,\text{GeV},\,M_{N_{2}}=347\,\text{GeV},\,{Y^{\prime}}_{11}=Y^{\prime}_{22}=1.2\times 10^{-9}\leavevmode\nobreak\ \text{GeV}^{-1}$
Figure 4: Evolution of the $\chi$ and $N(=N_{1}+N_{2})$ abundances in the early universe for the benchmark point (a) BP1 and (b) BP2. The observed DM relic density ($\Omega_{\chi}h^{2}\simeq 0.12$) is satisfied for both benchmark points.
Figure 5: The ratio of interaction rates to the Hubble rate $\Gamma/H$ for various processes as a function of $x=M_{\chi}/T$ for the benchmark point (a) BP1 and (b) BP2. The horizontal light blue line depicts $\Gamma/H$=1.
Fig. 4 shows the evolution of the densities of $\chi$ for the two benchmark points. Fig. 4 and  4 illustrate that the DM can chemically decouple at early times ($2\lesssim x\lesssim 4$) and still freeze-out at the late epoch of the Universe. On the contrary, $N_{1,2}$ stays in chemical equilibrium with the thermal bath until the freeze-out occurs. Note that $N_{1}$ rapidly decays to the DM after freeze-out. However, post freeze-out of $N_{1}$, since $Y_{N_{1}}<Y_{\chi}$, the decay term gives only a negligible contribution to the DM relic density.
The departure of $Y_{\chi}$ from chemical equilibrium at an early epoch, as shown in Fig. 4 and Fig. 4, arises due to $\chi$ pair annihilation, co-annihilation and $2\to 2$ inelastic processes going out of chemical equilibrium at the early epoch of the Universe. To illustrate this in more detail, we show in Fig 5 how $\Gamma/H$ varies as a function of $x$ for different processes for both benchmark points. We can see that all the $2\to 2$ processes involving $\chi$ decouple at high temperature (around $x\simeq 4$ for BP1 and $x\simeq 2.5$ for BP2 ). Consequently, $\chi$ undergoes early chemical decoupling in both BP1 and BP2 and ceases to track its equilibrium number density. Moreover, the inverse decay processes $\chi\nu_{e,\mu}\to N_{1,2}$ rate satisfies the following condition Frumkin:2021zng ; Frumkin:2025dxq ,
$\Gamma_{N_{1,2}\to\chi\nu_{e,\mu}}\lesssim\left(\frac{M_{N_{1,2}}-M_{\chi}}{M_{\chi}}\right)H(x=1),$ (13)
where $\Gamma_{N_{1,2}\to\chi\nu_{e,\mu}}$ is the partial decay width of $N_{1,2}$ to $\chi$. It ensures that inverse decay processes are too inefficient to restore chemical equilibrium, resulting in $Y_{\chi}>Y^{eq}_{\chi}$, note that $Y_{\chi}$ remains close to the equilibrium abundance. Furthermore, for BP1, $\chi\nu_{e,\mu}\to N_{1,2}$ both contribute significantly for $x\gtrsim 2$ which can be seen from the inset of Fig 5 thus ensuring DM freeze-out is primarily governed by $N_{1}$ freeze-out rather than by inverse decay freeze-out. In contrast, for BP2, the DM decouples when the inverse decay $\chi\nu_{e}\to N_{1}$ ceases.
It is also essential to bring to the attention of the readers that we have streamlined the calculation by assuming that $\chi$ remains in kinetic equilibrium, thus we solved the fully integrated Boltzmann equations. However, it was argued by the authors in Ref. Garny:2017rxs that dark matter may fail to maintain kinetic equilibrium along the process of conversion-driven freeze-out. A more accurate estimate of the relic abundance of the DM in the co-scattering mechanism is obtained through solving the coupled un-integrated Boltzmann equation. Nevertheless, since for our analysis the inverse decay process dominates over the $2\leftrightarrow 2$ processes, the result obtained using fully integrated Boltzmann equations are expected to differ mildly($\sim 10\%$) from those of the coupled un-integrated Boltzmann equation Brummer:2019inq .
4 Collider Constraints
Figure 6: (a) Branching ratio of $N_{2}$ vs its mass. (b) Production cross section of $N_{1}N_{2}$ at $\sqrt{s}=13$ TeV LHC vs $M_{N_{2}}$.
Figure 7: Feynman diagrams for the production of $N_{1}N_{2}$ at $p$ collider.
In this section we discuss how the LHC can constrain the parameter space of the model. The singlet fermions can be produced via the d=5 operators, ${c_{3}^{\prime}}\overline{N^{c}}\sigma_{\mu\nu}NB^{\mu\nu}$ with a sizable rate at the LHC. The heavier fermion $N_{2}$ decays to $N_{1}\gamma/N_{1}Z$ via the dipole operator or to $\chi\nu$ via Yukawa interactions.
Figure 8: Distribution of the photon $p_{T}$ for $M_{N_{2}}=[120,150,200]$ GeV and $M_{N_{1}}=105$ GeV.
Fig. 6 shows the branching ratio of $N_{2}$ as function of its mass $M_{N_{2}}$ for $M_{N_{1}},M_{\chi}=[105,100]\leavevmode\nobreak\ \rm{GeV}$, $c_{3}^{\prime}=Y_{22}^{\prime}=10^{-3}\leavevmode\nobreak\ \rm{GeV}^{-1}$. Note that there is no role of the coupling $Y_{11}^{\prime}$ in the decay of $N_{2}$. The decay $N_{2}\to N_{1}\gamma$ is the dominant channel as soon as it is kinematically accessible while $\text{Br}(N_{2}\to\chi\nu)\propto 1/{M_{N_{2}}^{2}}$ is strongly suppressed as $M_{N_{2}}$ increases. Moreover $Br(N_{2}\to N_{1}Z$) is subdominant and quickly saturates to $20\%$. For the choice of the $c_{3}^{\prime},Y_{22}^{\prime}$ to be considered (in in Table. 4)for the DM discussion, $\text{Br}(N_{2}\to\chi\nu)$ is extremely suppressed. Therefore, primary decay modes of $N_{2}$ will be $N_{2}\to N_{1}\gamma/N_{1}Z$. In Fig. 6, we show the production rate of $N_{1},N_{2}$ at the LHC. Feynman diagrams for the production processes are shown in Fig. 7. The dipole operator contributes to $pp\to N_{1}N_{2}$ and for this dominant production process the cross section scales as $(c_{3}^{\prime})^{2}$. For $N_{1}$, the only decay mode is $\chi\nu$. Hence, the dominant signature of $N_{2}$ will be mono-$\gamma$ associated with large missing momentum while mono-$Z$ is sub-dominant. The energy of $\gamma/Z$ depend directly on the mass splitting between $N_{2}$ and $N_{1}$. When $M_{N_{2}}-M_{N_{1}}$ is small, so that the photon is very soft, the mono-jet signal from $N_{1}N_{2}$ associated production with a hard-jet can become important. However, the production cross section for this process drops sharply as we demand high $p_{T}(j)$. In Fig. 6 we compare the production rate of $pp\to N_{1}N_{2}j$ (green line) for $p_{T}(j)\geq 200$ GeV with the one for $pp\to N_{1}N_{2}$ assuming a mass splitting of 5 GeV between $N_{2}$ and $N_{1}$. Note that the mono-jet signal can arise from the process $pp\to N_{i}N_{i}j$ via the coupling $c_{2}^{\prime}$, here it is irrelevant since we set $c_{2}^{\prime}=0$. The production of a DM pair via the coupling with Higgs, $pp\to h\to\chi\chi j$ can also lead to a mono-jet signal but the cross section is suppressed by the coupling $\lambda$ considered, $\lambda<10^{-6}$. We obtain that the mono-jet cross section in our model (for the parameter space of interest) is one order of magnitude smaller than the observed cross section from ATLAS search ATLAS:2021kxv . Therefore mono-$\gamma$ gives the leading signal. We consider only this process in the following.
Limit from mono -$\gamma$ search:
There exist an ATLAS search for events with an energetic photon and large missing transverse momentum at the 13 TeV LHC with integrated luminosity of $139\leavevmode\nobreak\ \text{fb}^{-1}$ ATLAS:2020uiq . As mentioned above, the production of $N_{1}N_{2}$ via the channel $pp\to N_{1}N_{2}$ mediated by the dipole operator ${c_{3}^{\prime}}\overline{N^{c}}\sigma_{\mu\nu}NB^{\mu\nu}$ can lead to such signal for the mass hierarchy $M_{N_{2}}>M_{N_{1}}>M_{\chi}$ such that $N_{2}\to N_{1}\gamma$ and $N_{1}$ decay invisibly, $N_{1}\to\chi\nu$ leading to $pp\to N_{1}N_{2}\to\gamma+E_{T}^{\rm miss}$. The mass splitting $M_{N_{2}}-M_{N_{1}}$ is crucial for the sensitivity of the signal as it governs the energy of the photon. Figure 8 shows the distribution of $p_{T}^{\gamma}$ for $M_{N_{2}}=[120,150,200]$ GeV and $M_{N_{1}},M_{\chi}=[105,100]$ GeV. As expected, the photon is harder for larger $M_{N_{2}}-M_{N_{1}}$. The signal regions in the ATLAS search ATLAS:2020uiq demands $E_{T}^{\gamma}>150$ GeV and upper limits at $95\%$ CL on the visible cross section are set for 7 signal regions (SRs): 4 inclusive (SRI1–SRI4) and 3 exclusive (SRE1-SRE3). The SRs are defined corresponding to different $E_{T}^{\rm miss}$ ranges, as listed in Table 3.
SRs SRI1 SRI2 SRI3 SRI4 SRE1 SRE2 SRE3
$E_{T}^{\rm miss}$ [GeV] >200>200 >250>250 >300>300 >375>375 200250200-250 250300250-300 300375300-375
$(\sigma\times A\times\epsilon)_{95}^{\rm obs}$ [fb] 2.45 1.42 0.93 0.53 1.80 1.04 0.79
Table 3: Upper limits at $95\%$ CL on the visible cross section $(\sigma\times A\times\epsilon)_{95}^{\rm obs}$ of the signal $pp\to\gamma+E_{T}^{\rm miss}$ for various signal regions (SRs) ATLAS:2020uiq .
Figure 9: Limit from ATLAS search ATLAS:2020uiq in the $(c_{3}^{\prime})_{UL}$ vs $M_{N_{2}}$ plane for signal region SRE1 (red) and SRI4 (black) for (a) $M_{\chi}=100$ GeV, $M_{N_{1}}=105$ GeV (b) $M_{\chi}=300$ GeV, $M_{N_{1}}=315$ GeV. In both cases $Y_{22}^{\prime}=10^{-9}\leavevmode\nobreak\ {\rm GeV}^{-1}$.
Figure 10: Upper limit on $c_{3}^{\prime}$ from the ATLAS search ATLAS:2020uiq in the $\delta_{1}-\delta_{2}$ plane for (a) $M_{\chi}=100$ GeV and (b) $M_{\chi}=300$ GeV. Here $Y_{22}^{\prime}=10^{-9}\leavevmode\nobreak\ {\rm GeV^{-1}}$.
The Lagrangian of this model is implemented in FeynRules(v2.3) Alloul:2013bka and the generated UFO files are used in the MC event generator MADGRAPH5(v3.4) Alwall:2014hca to simulate the signal events. Partonic events are passed through PYTHIA8 Sjostrand:2014zea to perform showering and hadronization. We use Delphes (3.5) deFavereau:2013fsa for the simulation of detector effect and to calculate the signal efficiencies.
Following the selection cuts used in the ATLAS search ATLAS:2020uiq , we calculated the selection efficiencies for our signal model. We demand that the leading photon has $E_{T}^{\gamma}>150$ GeV, $1.52<|\eta^{\gamma}|<2.37$ or $|\eta^{\gamma}|<1.37$, and $\Delta\phi(\gamma,E_{T}^{\rm miss})>0.4$. Then we calculate the signal events in the seven SRs and set limit in the $M_{N_{2}}$ vs $(c_{3}^{\prime})_{UL}$ plane as shown in Fig. 9, where $(c_{3}^{\prime})_{UL}$ represent the $95\%$ CL upper limit on $c_{3}^{\prime}$. Here we only show the strongest and weakest limit. We also calculate the HL-LHC projection, for this we scale the observed upper limit on the visible cross section by a factor $\sqrt{139/3000}$ and compare with the theory prediction. Fig. 9 is for DM mass $M_{\chi}=100$ GeV, $M_{N_{1}}=105$ GeV and $Y_{22}^{\prime}=10^{-9}\leavevmode\nobreak\ \rm{GeV}^{-1}$. Here the solid lines indicate current LHC limit while dashed lines are for HL-LHC projection. The strongest limit is set in SRI4 (black) and the weakest limit in SRE1 (red). $(c_{3}^{\prime})_{UL}$ is determined by the production cross section $\sigma(pp\to N_{1}N_{2})$, $Br(N_{2}\to N_{1}\gamma)$ and cut efficiency $\epsilon$. For a fixed $M_{N_{1}}$ and $c_{3}^{\prime}$, the $\sigma(pp\to N_{1}N_{2})$ drops with $M_{N_{2}}$. While, $Br(N_{2}\to N_{1}\gamma)$ remains steady except near the kinematic threshold, as evident from Fig.6. The cut efficiency is different for different SRs. For SRI4, $\epsilon$ increases with $M_{N_{2}}$ when $M_{N_{1}}$ is fixed and it can vary from $10^{-4}$ for $M_{N_{2}}=110$ GeV, to $0.65$ for $M_{N_{2}}=1000$ GeV. Whereas, for SRE1, $\epsilon$ increases initially with $M_{N_{2}}$ and at a certain point it starts dropping. Therefor, $(c_{3}^{\prime})_{UL}$ varies differently for different SRs. The limit is weakest when $M_{N_{2}}\simeq M_{N_{1}}$ as the photon becomes very soft. For $M_{N_{2}}\simeq 200$ GeV limit on $c_{3}^{\prime}$ becomes strongest. In the region $M_{N_{2}}>200$ GeV, for SRI4, the rise in $\epsilon$ gets canceled by the drop in $\sigma$ leading to a almost flat limit on $c_{3}^{\prime}$ w.r.t $M_{N_{2}}$. Whereas for SRE1, the limit weakens as $\sigma$ drops faster compared to the variation in $\epsilon$. Fig. 9 shows the limits in the case $M_{\chi}=300$ GeV, $M_{N_{1}}=315$ GeV and $Y_{22}^{\prime}=10^{-9}\leavevmode\nobreak\ \rm{GeV}^{-1}$ . The dotted line represents $\Gamma(N_{2})/M_{N_{2}}=0.01$. Finally, in Fig. 10, we show the limit on $c_{3}^{\prime}$ in the $\delta_{1}$ vs $\delta_{2}$ plane, as defined in Eq. 11. Here we consider only SRI4 since it sets the strongest limit.
Figure 11: (a) DM relic density and (b) $\Delta^{1}_{\chi}$ in the $\delta_{1}$ and $\delta_{2}$ plane for $M_{\chi}=100\,\text{GeV}$, $c_{3}^{\prime}=7\times 10^{-4}\,\rm{GeV^{-1}},Y^{\prime}_{11,22}=1.1\times 10^{-9}\,\rm{GeV^{-1}}$. The dash- dotted line satisfies the DM observed relic density.
5 Numerical Analysis
As discussed in Section 3, we re-emphasize that the DM $\chi$ gets diluted in the early Universe either through co-annihilation, which is primarily governed by $N_{1,2}$ decoupling from the thermal bath or through the inelastic processes co-scattering. The inverse decay process primarily governs the inelastic processes, i.e. $\chi\nu_{e,\mu}\to N_{1,2}$. To delineate the parameter space of our model, we first concentrate on the two benchmark points, BP1 and BP2, introduced in Section 3 but with varying $M_{N_{1}}$ and $M_{N_{2}}$ and keeping other parameters fixed. This is followed by a comprehensive exploration of the model’s parameter space.
In Fig. 11, we show the variations of the DM relic density contours in the $\delta_{1}-\delta_{2}$ plane for BP1. The black dashed-dot line corresponds to $\Omega_{\chi}h^{2}=0.12$. When $\delta_{2}$ is less than $0.15$ the contour line corresponding to $\Omega_{\chi}h^{2}=0.12$ is approximately independent of $\delta_{2}$. It arises because of efficient dilution of $N_{1,2}$ through co-annihilation as $\delta_{2}$ decreases, and the relic density is set when $N_{1}$ decouples through the thermal bath. Figure 11 shows that $\Delta^{1}_{\chi}$ decreases as $\delta_{2}$ decreases. Hence, in this limit, the standard freeze-out via co-annihilation is realized. On the contrary, for $\delta_{1}$ less than $5\times 10^{-2}$, the inverse decay rate is suppressed due to phase space suppression and satisfies the condition given in Eq. 13. Furthermore, the dilution of $N_{1,2}$ is not efficient owing to the large mass splitting between $N_{1}$ and $N_{2}$. Hence, the relic density gets frozen when the inverse decay of $\chi$ to $N_{1}$ ceases. It is in the spirit of a co-scattering mechanism, and $\Delta^{1}_{\chi}$ approaches 1. Also, note that DM is overabundant for large $\delta_{1}$ and $\delta_{2}$ due to the fact that neither the inverse decay nor (co-)annihilation of $N_{1,2}$ are efficient. The black dashed and dotted line show the current LHC limit and HL-LHC future projection in the $\delta_{1}-\delta_{2}$ plane. As illustrated in Fig. 11, the co-scattering dominated region is subject to stringent constraints from LHC ATLAS:2021kxv , with only values of $\Delta^{1}_{\chi}\lesssim 0.4$ allowed when imposing that $\Omega h^{2}=0.12$. The HL-LHC has the potential to probe the co-scattering region when $\Delta^{1}_{\chi}>0.2$ .
Figure 12: Same as Fig. 11 for $M_{\chi}=300\,\text{GeV}$.
Parameter  Scanned range
$M_{\chi}\,\,\rm{[GeV]}$  [$100$ , $500$]
$\delta_{1}(M_{N_{1}}\leavevmode\nobreak\ \rm{[GeV]})$  $10^{-3}$ , $0.5$] $([\approx M_{\chi},750])$
$\delta_{2}(M_{N_{2}}\leavevmode\nobreak\ \rm{[GeV]})$  $0^{-3} , 11$] $([\approx M_{N_{1}},1500])$
$Y^{\prime}_{11}\,\,\rm{[GeV^{-1}]}$  $10^{-11}$ , $10^{-6}$]
$Y^{\prime}_{22}\,\,\rm{[GeV^{-1}]}$  $10^{-11}$ , $10^{-6}$]
$c^{\prime}_{3}\,\,\rm{[GeV^{-1}]}$  $10^{-6}$ , $10^{-3}$]
Table 4: Input parameters used in the numerical scan to determine the allowed parameter space. Note that we have assumed $c^{\prime}_{1,2}=0$ and $\lambda=10^{-6}$.
Figure 13: Fig. 13 illustrates (dis)allowed parameter space in $M_{N_{1}}-M_{N_{2}}$ plane, subject to relic density constraints $10^{-4}\leq\Omega_{\chi}h^{2}\leq 0.1224$ along with collider limit. The color pallet corresponds to $c^{\prime}_{3}$. Fig. 13 shows variations of $\Omega_{\chi}h^{2}$ vs $c^{\prime}_{3}$ with the DM mass $M_{\chi}$ in color pallet after imposing the LHC constraint. Fig. 13 shows variations of $Y^{\prime}_{11}$ vs $Y^{\prime}_{22}$ with the $\Delta^{1}_{\chi}$ in the color pallet. Note that we have assumed $c^{\prime}_{1,2}=0$ and $\lambda=10^{-6}$. Fig. 13 Shows parameter space that can be probed at the HL-LHC.
Following the same approach as above, we show in Fig. 12 the contours of DM $\chi$ relic density and $\Delta^{1}_{\chi}$ for $M_{\chi}=300\,\rm{GeV}$. As before, we have fixed $c^{\prime}_{3}=7\times 10^{-4}\,\rm{GeV^{-1}}$. However, the allowed masses for $N_{1}$ and $N_{2}$ in BP2 are larger than those in BP1 to ensure that $\chi$ be the DM candidate. This results in a decrease in $N_{1,2}$ (co-)annihilation cross-section for BP2 compared to BP1 and, in turn, facilitates early $N_{1,2}$ decoupling from the thermal bath in BP2. Therefore, the relic density of $\chi$ is primarily governed by the inverse decay process of $\chi$ to $N_{1,2}$ and results in a large value of $\Delta^{1}_{\chi}$, see Fig. 12. In this case, the LHC does not constrain the parameter space where the relic density matches the observed value as it requires a too small value of $\delta_{2}$, as seen from the black dashed and dot-dashed lines in Fig. 12. However the HL-LHC can probe some of the parameter space in the co-scattering dominant region.
We comprehensively explored the model’s parameter space through a flat random scan over the free parameters within the ranges shown in Table 4. We consider the points which satisfy the DM relic density in the range,
$10^{-4}\leq\Omega_{\chi}h^{2}\leq 0.1224.$ (14)
In Fig. 13, we show the allowed range of $M_{N_{1}}$ vs $M_{N_{2}}$ with $c^{\prime}_{3}$ in the color pallet after imposing the relic density constraint as given by Eq. 14. We then impose the ATLAS mono-$\gamma$ limit as discussed in Sec. 4. Figure  13 shows that the mono-$\gamma$ search strongly constrain scenarios with large mass splitting between $M_{N_{1}}$ and $M_{N_{2}}$, these points are shown in red in the figure. Furthermore, as discussed above, the relic density constraint requires at least $N_{1}$ to be present in the early Universe, so that $\chi$ can be diluted through co-annihilation or inverse decay. Thus the mass difference between $M_{N_{1}}$ and $M_{\chi}$ cannot be too large, and we obtain an upper limit on $M_{N_{1}}<600\,\rm{GeV}$. It is important to notice that $N_{2}$ can also participate in DM $\chi$ dilution through co-annihilation or inverse decay. However since the contribution of $N_{1}$ is always present, the upper bound on $M_{N_{2}}$ is less stringent, specifically, $M_{N_{2}}\leq 1000\,\mathrm{GeV}$.
In Fig. 13, we depict the allowed points after the relic density constraint and collider limit in the $\Omega_{\chi}h^{2}$ and $c^{\prime}_{3}$ plane with the color pallet corresponding to the DM mass $M_{\chi}$. The points overlapping with the red line satisfy the observed DM relic density within the $2\sigma$ range. As $c^{\prime}_{3}$ decreases, $N_{1,2}$ undergoes early chemical decoupling from the thermal bath, which in turn does not contribute in $\chi$ dilution neither through co-annihilation nor through co-scattering. We can see that for the range of $M_{\chi}$ that survived $c^{\prime}_{3}\leq 3\times 10^{-4}$ , i.e. $M_{\chi}>200\,\rm{GeV}$, corresponds to $M_{N_{1,2}}>M_{\chi}$. It happens because, for $M_{N_{1,2}}>200\,\rm{GeV}$ additional channels such as $N_{1}N_{2}\to t\bar{t}$ become kinematically accessible. In Fig. 13, we show the variation of the allowed points in $Y^{\prime}_{11}$ - $Y^{\prime}_{22}$ plane with the color pallet corresponding to $\Delta^{\chi}_{1}$. As evident from Fig. 13, for large $Y^{\prime}_{11,22}$, $\chi$ thermalizes with the thermal bath due to rapid decay and inverse decay as $\Delta^{\chi}_{1}\to 0$, and the relic density is governed by the thermal decoupling of $N_{1,2}$. It is the region of parameter space where a co-annihilation process governs $\Omega_{\chi}h^{2}$. On the contrary, for $Y^{\prime}_{11,22}<10^{-8}\,\rm{GeV^{-1}}$, $\chi$ undergoes early chemical decoupling and $\Omega_{\chi}h^{2}$ gets frozen when $\chi$ inverse decay to $N_{1,2}$ ceases. Consequently, in this region of parameter space, the value of $\Delta^{\chi}_{1}>0.5$, indicating that the DM relic density is predominantly determined by the co-scattering mechanism. In Fig 13, we show the discovery prospects via the mono-$\gamma$ signal at the HL-LHC. The dotted-color points in the $M_{N_{1}}$ and $M_{N_{2}}$ plane can be probed in HL-LHC while the red-crossed points remain out of HL-LHC reach. These points correspond to lower values of $M_{N_{2}}-M_{N_{1}}$ and/or lower value of $c^{\prime}_{3}$, which leads to suppressed signal efficiency and/or suppressed production cross section for the signal $pp\to N_{2}N_{2}\to\gamma+E_{T}^{\rm miss}$.
6 Summary and Conclusion
In this work, we investigate the dark matter production mechanism in an EFT framework where we extent the SM with two singlet fermions $N_{1,2}$ and a real singlet scalar $\chi$, which are $Z_{2}$ odd particles. The singlet scalar $\chi$ serves as a dark matter candidate. Both $\chi$ and $N_{1,2}$ are thermal particles and their interactions in the early Universe controls the DM abundance. We emphasize the role of dimension-5 operators in the co-scattering mechanism. The key parameters on which DM dynamics depend are the masses $M_{N_{1,2}},M_{\chi}$ and couplings $\lambda,c_{3}^{\prime},Y_{11,22}$. Note that, we have assumed $c_{1,2}^{\prime}=0$ throughout our analysis.
We first discuss DM formation using two benchmarks, BP1 where the co-annihilation mechanism dominates and BP2 where co-scattering dominates. We fix the Higgs portal coupling $\lambda=10^{-6}$ ensuring that $\chi$ remains in thermal equilibrium and also to evade stringent direct detection constraints. The DM relic density in BP1 is set when $N_{1}$ decouples from the thermal bath and the abundance is determined by the freeze-out of processes that maintain equilibrium between states until decoupling. In contrast, for BP2, the relic density is set when the inverse decay process ceases and is governed by the co-scattering mechanism, where conversions between states—rather than annihilations—control the final abundance. In both cases we find that the inverse decay rates, $\chi,\nu_{e,\mu}\to N_{1,2}$, although more efficient in BP1, are not sufficient to bring $\chi$ into chemical equilibrium with the thermal bath. We also find as expected that small mass splittings ($\delta_{1}<0.1$ or $\delta_{2}<0.1$) are needed to prevent overproduction of DM.
As concerns the collider constraints, we find that the dimension-5 operators lead to sizable production of $N_{1,2}$ at the LHC and that the most stringent constraint is coming from the mono-$\gamma$ signature via $pp\to N_{1}N_{2}(\to N_{1}\gamma)\to\gamma+E_{T}^{\rm miss}$. Following the ATLAS analysis, we calculate the sensitivity of this signal which primarily depends on the mass difference $M_{N_{2}}-M_{N_{1}}=\delta_{2}$ (or the photon energy) and $c_{3}^{\prime}$. We calculate the mono-$\gamma$ limit on $c_{3}^{\prime}$ in the $\delta_{2}-\delta_{1}$ plane, where $\delta_{1}$ dictates the relative mass splitting between $M_{N_{1}}$ and $M_{\chi}$. The stronger bound on $c_{3}^{\prime}$ is found for larger $\delta_{2}$ since the energy of the photon grows with $\delta_{2}$.
[…]
In conclusion, we have investigated the production of dark matter, with a primary focus on the co-scattering mechanism within the framework of EFT. This framework offers rich dark matter phenomenology that can be comprehensively probed at the HL-LHC. While we concentrate on the dipole operator, other d=5 operators could play a crucial role for dark matter production, these will be studied in future work.
Acknowledgements
The work of AR is supported by Basic Science Research Program through the National Research Foundation of Korea(NRF) funded by the Ministry of Education through the Center for Quantum Spacetime (CQUeST) of Sogang University (RS-2020-NR049598). RP is supported in part by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education, Science and Technology (NRF-2022R1A2C2003567). GB thanks IFIRSE, Quy Nhon, Vietnam and Rencontres du Vietnam for their warm hospitality while this work was finalised. This work was funded in part by the Indo-French Centre for the Promotion of Advanced Research (Project title: Beyond Standard Model Physics with Neutrino and Dark Matter at Energy, Intensity and Cosmic Frontiers, Grant no: 6304-2). The authors also acknowledges SAMKHYA: High-Performance Computing Facility provided by the Institute of Physics (IoP), Bhubaneswar.

### v2

@@
 $+\lambda\Phi^{\dagger}\Phi\chi^{2}+\beta\chi^{4}.$
-In the above $\tilde{\Phi}=i\sigma_{2}\Phi^{\star}$ and $N^{c}=(i\gamma^{2}\gamma^{0})\overline{N}^{T}$, where $N=(N_{1},N_{2})^{T}$.
+In the above equation $L$, $\Phi$ and $B_{\mu\nu}$ are the SM Lepton doublet, Higgs doublet and hypercharge field strength tensor, respectively; additionally, $\tilde{\Phi}=i\sigma_{2}\Phi^{\star}$ and $N^{c}=(i\gamma^{2}\gamma^{0})\overline{N}^{T}$, where $N=(N_{1},N_{2})^{T}$.
 Note that in the above Lagrangian flavor indices are implicit.
@@
 Therefore, the field $N_{i}$ can not participate in the SM neutrino mass generation at tree level via the d=5 Weinberg operator and there is no mixing between SM neutrino and $N_{i}$.
-After EW symmetry breaking, $c_{2}$ contribute to the mass of $N$ as $M_{N}=M_{B}+c_{2}v^{2}/\Lambda$.
+After electroweak symmetry breaking, $c_{2}$ contribute to the mass of $N$ as $M_{N}=M_{B}+c_{2}v^{2}/\Lambda$.
 We redefine the couplings of the d=5 operators as, $c_{5}^{\prime}=\frac{c_{5}}{\Lambda},y^{\prime}=\frac{y}{\Lambda},c_{1}^{\prime}=\frac{c_{1}}{\Lambda},c_{2}^{\prime}=\frac{c_{2}}{\Lambda},c_{3}^{\prime}=\frac{c_{3}}{\Lambda}$.
@@
 We consider both $\chi$ and $N_{1,2}$ to be in thermal equilibrium with the SM plasma. The thermal equilibrium of $\chi$ is maintained by the following processes,
-  • $\chi$ pair annihilation: $\chi\chi\to\text{SM}\leavevmode\nobreak\ \text{SM}$.
-  • $\chi$ and $N_{1,2}$ co-annihilation: $\chi\leavevmode\nobreak\ N_{1,2}\to\text{SM}\leavevmode\nobreak\ \text{SM}$.
-  • $N_{1,2}/\chi$ interchange: $N_{1,2}\leavevmode\nobreak\ \text{SM}\to\chi\leavevmode\nobreak\ \text{SM}$, $\chi\leavevmode\nobreak\ \nu_{e,\mu}\to N_{1,2}$ .
+  • $\chi$ pair annihilation: $\chi\chi\to\text{SM}\penalty 10000\ \text{SM}$.
+  • $\chi$ and $N_{1,2}$ co-annihilation: $\chi\penalty 10000\ N_{1,2}\to\text{SM}\penalty 10000\ \text{SM}$.
+  • $N_{1,2}/\chi$ interchange: $N_{1,2}\penalty 10000\ \text{SM}\to\chi\penalty 10000\ \text{SM}$, $\chi\penalty 10000\ \nu_{e,\mu}\to N_{1,2}$ .
 Furthermore, the thermal equilibrium of $N_{1,2}$ is maintained either by the same interactions or by the following processes,
-  • $N_{1,2}$ pair annihilation: $N_{1,2}\leavevmode\nobreak\ N_{1,2}\to\gamma\gamma,\gamma Z,ZZ$.
-  • $N_{1}$ and $N_{2}$ co-annihilation: $N_{1}\leavevmode\nobreak\ N_{2}\to f\bar{f},W^{+}W^{-},ZH$.
+  • $N_{1,2}$ pair annihilation: $N_{1,2}\penalty 10000\ N_{1,2}\to\gamma\gamma,\gamma Z,ZZ$.
+  • $N_{1}$ and $N_{2}$ co-annihilation: $N_{1}\penalty 10000\ N_{2}\to f\bar{f},W^{+}W^{-},ZH$.
 We assume $\chi$ to be the DM particle and $N_{1,2}$ the next-to-lightest-odd (NLOP) particles.
@@
 1, in addition $\chi\nu_{e,\mu}\to N_{1,2}$ contribute to co-scattering.
-For convenience and to keep with the formalism used in micrOMEGAs Alguero:2022inz we separate all particles into three sectors, namely, sector 0 corresponding to SM particles, sector $1$ comprising $N_{1}$ and $N_{2}$, and sector $2$ containing $\chi$.
+For convenience and to keep with the formalism used in micrOMEGAs Alguero:2022inz we separate all particles into three sectors, namely, sector 0 corresponding to SM particles, sector $1$ comprising $N_{1}$ and $N_{2}$, and sector $2$ containing $\chi$  [^1] Here we will refer to the dark sectors either as sector 1 and sector 2 or as $N\,(=N_{1}+N_{2})$ and $\chi$, respectively.
+.
 The particles within each sector are assumed to be in thermal equilibrium.
@@
 Initial state Final state Scaling with couplings
-$\chi$ $h$ $N_{1,2}$ $\nu_{e,\mu}$ ${Y^{{}^{\prime}}}_{11(22)}^{2}$
-$\chi$ $W^{\pm}$ $N_{1,2}$ $e^{\pm}\mu^{\pm}$ ${Y^{{}^{\prime}}}_{11(22)}^{2}$ (t-channel process )
-$\chi$ $\nu_{e,\mu}$ $N_{1,2}$ - ${Y^{{}^{\prime}}}_{11(22)}^{2}$ (Inverse Decay)
+$\chi$ $h$ $N_{1,2}$ $\nu_{e,\mu}$ ${y^{\prime}}_{11(22)}^{2}$
+$\chi$ $W^{\pm}$ $N_{1,2}$ $e^{\pm}\mu^{\pm}$ ${y^{\prime}}_{11(22)}^{2}$ (t-channel process )
+$\chi$ $\nu_{e,\mu}$ $N_{1,2}$ - ${y^{\prime}}_{11(22)}^{2}$ (Inverse Decay)
 Table 1: List of all relevant conversion processes shown in Fig. 2. All cross sections depend on the square of the Yukawa coupling $y^{\prime}$.
@@
 $\Delta^{1}_{\chi}\equiv 1-\frac{\Omega h^{2}(\textrm{Single})}{\Omega h^{2}(\textrm{Coupled})}.$ (8)
-More specifically, $\Omega h^{2}(\text{Single})$ refers to the value obtained using the standard darkOmega function in micrOMEGAs, which accounts for only one Boltzmann equation and thus incorporates solely co-annihilation processes.
-In contrast, $\Omega h^{2}(\text{Coupled})$ represents the result obtained via the darkOmegaN function, where the dark particles are divided into two sectors.
+More specifically, $\Omega h^{2}(\text{Single})$ refers to the value obtained using the standard darkOmega function in micrOMEGAs, which accounts for only one Boltzmann equation and assumes that all particles are in thermal equilibrium in the early Universe.
+Thus it incorporates solely co-annihilation processes.
+In contrast, $\Omega h^{2}(\text{Coupled})$ represents the result obtained via the darkOmegaN function, where the dark particles are divided into two sectors that are not necessarily in thermal equilibrium with each other.
 This calculation involves solving the coupled system of Boltzmann equations 2 and 3, including all co-scattering and decay processes.
 Figure 3: Feynman diagrams for the pair annihilation and co-annihilation process of $N_{1,2}$ that depend on the coupling $c_{3}^{\prime}$.
-Note that due to the choice $c^{\prime ij}_{3}=\epsilon^{ij}c^{\prime ij}_{3}$, we have both co-annihilation and annihilation processes between $N_{1}$ and $N_{2}$.
+Note that due to the choice $(c^{\prime}_{3})_{ij}=\epsilon_{ij}(c^{\prime}_{3})_{ij}$, we have both co-annihilation and annihilation processes between $N_{1}$ and $N_{2}$.
 Initial state Final state Scaling with couplings
-$N_{1,2}$ $N_{1,2}$ $\gamma,Z$ $\gamma,Z$ ${c_{3}^{{}^{\prime}}}^{4}$ (t- channel process )
-$N_{1}$ $N_{2}$ $f$ $\bar{f}$ ${c_{3}^{{}^{\prime}}}^{2}$ (s-channel process )
-$N_{1}$ $N_{2}$ $^{+} WW^{-}$ ${c_{3}^{{}^{\prime}}}^{2}$ (s-channel process )
-$N_{1}$ $N_{2}$ $Z$ $H$ ${c_{3}^{{}^{\prime}}}^{2}$ (s-channel process )
-Table 2: List of all relevant annihilation and co-annihilation processes as in Fig. 3, with the dependency of their cross sections on the $c_{3}^{\prime}=c_{3}^{{}^{\prime}12}$.
-Note that both sectors remain in thermal equilibrium when the Wilson coefficient ($c_{i}^{{}^{\prime}},\,Y^{{}^{\prime}}$) and the Higgs portal coupling $\lambda$ are sufficiently large.
+$N_{1,2}$ $N_{1,2}$ $\gamma,Z$ $\gamma,Z$ ${c_{3}^{\prime}}^{4}$ (t- channel process )
+$N_{1}$ $N_{2}$ $f$ $\bar{f}$ ${c_{3}^{\prime}}^{2}$ (s-channel process )
+$N_{1}$ $N_{2}$ $^{+} WW^{-}$ ${c_{3}^{\prime}}^{2}$ (s-channel process )
+$N_{1}$ $N_{2}$ $Z$ $H$ ${c_{3}^{\prime}}^{2}$ (s-channel process )
+Table 2: List of all relevant annihilation and co-annihilation processes as in Fig. 3, with the dependency of their cross sections on the $c_{3}^{\prime}=(c_{3}^{\prime})_{12}$.
+Note that both sectors remain in thermal equilibrium when the Wilson coefficient ($c_{i}^{\prime},\,y^{\prime}$) and the Higgs portal coupling $\lambda$ are sufficiently large.
 If this is the case $Y_{1}/Y_{2}=Y_{1}^{eq}/Y_{2}^{eq}$ and Eq.
-23 simplify considerably and solving the coupled Boltzmann equations yields the same result as solving a single abundance equation where $Y_{1}=0$ and $Y_{2}=Y$.
+2, 3 simplify considerably and solving the coupled Boltzmann equations yields the same result as solving a single abundance equation where $Y_{1}=0$ and $Y_{2}=Y$.
 Hence, the usual freeze-out result is recovered and $\Delta^{1}_{\chi}=0$.
-Conversely, in scenarios where $Y^{{}^{\prime}},\lambda\sim\mathcal{O}(10^{-6}-10^{-10})$, $\chi$’s pair annihilation, co-annihilation, and exchange process involving $\chi$ and constituents of sector 2 become negligible, thereby simplifying the Boltzmann equations [^1] For the numerical analysis, we solve the general Boltzmann equations governing the evolution of $\chi$ and the constituents of sector 1, specifically Eqs.
+Conversely, in scenarios where $y^{\prime},\lambda\sim\mathcal{O}(10^{-6}-10^{-10})$, $\chi$’s pair annihilation, co-annihilation, and exchange process involving $\chi$ and constituents of sector 2 become negligible, thereby simplifying the Boltzmann equations [^2] For the numerical analysis, we solve the general Boltzmann equations governing the evolution of $\chi$ and the constituents of sector 1, specifically Eqs.
 2 and 3.,
@@
 2.
-All cross-sections depend on ${Y^{{}^{\prime}}}^{2}$ as shown in Table 1.
+All cross-sections depend on $y^{\prime 2}$ as shown in Table 1.
 It is important to emphasize that we need at least $N_{1}$ to be in chemical equilibrium with the thermal bath through (co-)annihilation process, i.e.
 1100, which depends on $c_{3}^{\prime}$ solely, to facilitate $\chi$ dilution either through co-annihilation or co-scattering.
+Note that 1100 refers to processes of the type $N_{i}N_{j}\to SM,SM$, these are illustrated in Fig.
+3.
 Therefore, it becomes necessary to assume $c_{3}^{\prime}>10^{-5}$ to ensure that the constituents of sector 1 remain in chemical equilibrium at $T\geq M_{N_{1,2}}$.
@@
 The relevant Feynman diagrams for $N_{1}$ dilution in the early Universe are shown in Fig.
-3, and the interaction rates dependence on $c_{3}^{{}^{\prime}}$ are given in Table 2.
+3, and the interaction rates dependence on $c_{3}^{\prime}$ are given in Table 2.
+It is important to bring to the attention of the readers that a wide range of DM masses is feasible in our model, spanning from sub-GeV to the TeV scale.
+In our work, we primarily focus on weak-scale DM mass to have a viable collider signature in the LHC, as will be discussed in the next section.
+In contrast, neither upcoming direct nor indirect detection experiments can provide viable signatures, since the coupling responsible for dark matter detection, $\lambda$, is set to $10^{-6}$.
 We begin by examining two benchmark points to illustrate how the evolution of DM $\chi$ depends on the key parameters, i.e.
@@
 Based on the primary production mechanism of $\chi$, either through co-annihilation or co-scattering, we have considered two benchmark points.
-We assume $c_{3}^{\prime}=7\times 10^{-4}\leavevmode\nobreak\ \text{GeV}^{-1}$ and the other parameters are set to,
-  • BP1: $M_{\chi}=100\,\text{GeV},\,M_{N_{1}}=105\,\text{GeV},\,M_{N_{2}}=115\,\text{GeV},\,{Y^{\prime}}_{11}=Y^{\prime}_{22}=1.1\times 10^{-9}\leavevmode\nobreak\ \text{GeV}^{-1}$
-  • BP2: $M_{\chi}=300\,\text{GeV},\,M_{N_{1}}=315\,\text{GeV},\,M_{N_{2}}=347\,\text{GeV},\,{Y^{\prime}}_{11}=Y^{\prime}_{22}=1.2\times 10^{-9}\leavevmode\nobreak\ \text{GeV}^{-1}$
+We assume $c_{3}^{\prime}=7\times 10^{-4}\penalty 10000\ \text{GeV}^{-1}$ and the other parameters are set to,
+  • BP1: $M_{\chi}=100\,\text{GeV},\,M_{N_{1}}=105\,\text{GeV},\,M_{N_{2}}=115\,\text{GeV},\,{y^{\prime}}_{11}=y^{\prime}_{22}=1.1\times 10^{-9}\penalty 10000\ \text{GeV}^{-1}$
+  • BP2: $M_{\chi}=300\,\text{GeV},\,M_{N_{1}}=315\,\text{GeV},\,M_{N_{2}}=347\,\text{GeV},\,{y^{\prime}}_{11}=y^{\prime}_{22}=1.2\times 10^{-9}\penalty 10000\ \text{GeV}^{-1}$
 Figure 4: Evolution of the $\chi$ and $N(=N_{1}+N_{2})$ abundances in the early universe for the benchmark point (a) BP1 and (b) BP2.
@@
 In contrast, for BP2, the DM decouples when the inverse decay $\chi\nu_{e}\to N_{1}$ ceases.
-It is also essential to bring to the attention of the readers that we have streamlined the calculation by assuming that $\chi$ remains in kinetic equilibrium, thus we solved the fully integrated Boltzmann equations.
-However, it was argued by the authors in Ref.
-Garny:2017rxs that dark matter may fail to maintain kinetic equilibrium along the process of conversion-driven freeze-out.
-A more accurate estimate of the relic abundance of the DM in the co-scattering mechanism is obtained through solving the coupled un-integrated Boltzmann equation.
-Nevertheless, since for our analysis the inverse decay process dominates over the $2\leftrightarrow 2$ processes, the result obtained using fully integrated Boltzmann equations are expected to differ mildly($\sim 10\%$) from those of the coupled un-integrated Boltzmann equation Brummer:2019inq .
+It is important to point out that we have streamlined the calculation by assuming that $\chi$ remains in kinetic equilibrium, thus we solved the fully integrated Boltzmann equations.
+In our analysis, the inverse decay process dominates over the $2\leftrightarrow 2$ interactions.
+Consequently, the results obtained using the fully integrated Boltzmann equations are expected to differ by less than $10\%$ from those of the coupled un-integrated Boltzmann equations Garny:2017rxs ; Alguero:2022inz .
+This small deviation arises because the decay and inverse decay processes remain efficient until late times, thereby re-establishing an approximate kinetic equilibrium (see Fig.
+8 of Ref.
+Garny:2017rxs ).
+However, in scenarios where the inverse decay becomes subdominant, the deviation from the coupled un-integrated Boltzmann results can become significant, as the dark matter momentum distribution becomes highly distorted (i.e., deviates considerably from a thermal distribution).
+In such cases, the integrated Boltzmann equation may under- or overestimate the actual relic yield Brummer:2019inq , thus solving the coupled un-integrated Boltzmann equation will provide a more accurate estimate of the relic abundance of the DM.
+We leave this for future work.
 4 Collider Constraints
@@
 Fig.
-6 shows the branching ratio of $N_{2}$ as function of its mass $M_{N_{2}}$ for $M_{N_{1}},M_{\chi}=[105,100]\leavevmode\nobreak\ \rm{GeV}$, $c_{3}^{\prime}=Y_{22}^{\prime}=10^{-3}\leavevmode\nobreak\ \rm{GeV}^{-1}$.
+6 shows the branching ratio of $N_{2}$ as function of its mass $M_{N_{2}}$ for $M_{N_{1}},M_{\chi}=[105,100]\penalty 10000\ \rm{GeV}$, $c_{3}^{\prime}=y_{22}^{\prime}=10^{-3}\penalty 10000\ \rm{GeV}^{-1}$.
 Note that there is no role of the coupling $Y_{11}^{\prime}$ in the decay of $N_{2}$.
@@
 Moreover $Br(N_{2}\to N_{1}Z$) is subdominant and quickly saturates to $20\%$.
-For the choice of the $c_{3}^{\prime},Y_{22}^{\prime}$ to be considered (in in Table.
+For the choice of the $c_{3}^{\prime},y_{22}^{\prime}$ to be considered (in Table.
 4)for the DM discussion, $\text{Br}(N_{2}\to\chi\nu)$ is extremely suppressed.
@@
 Limit from mono -$\gamma$ search:
-There exist an ATLAS search for events with an energetic photon and large missing transverse momentum at the 13 TeV LHC with integrated luminosity of $139\leavevmode\nobreak\ \text{fb}^{-1}$ ATLAS:2020uiq .
+There exist an ATLAS search for events with an energetic photon and large missing transverse momentum at the 13 TeV LHC with integrated luminosity of $139\penalty 10000\ \text{fb}^{-1}$ ATLAS:2020uiq .
 As mentioned above, the production of $N_{1}N_{2}$ via the channel $pp\to N_{1}N_{2}$ mediated by the dipole operator ${c_{3}^{\prime}}\overline{N^{c}}\sigma_{\mu\nu}NB^{\mu\nu}$ can lead to such signal for the mass hierarchy $M_{N_{2}}>M_{N_{1}}>M_{\chi}$ such that $N_{2}\to N_{1}\gamma$ and $N_{1}$ decay invisibly, $N_{1}\to\chi\nu$ leading to $pp\to N_{1}N_{2}\to\gamma+E_{T}^{\rm miss}$.
@@
 $(\sigma\times A\times\epsilon)_{95}^{\rm obs}$ [fb] 2.45 1.42 0.93 0.53 1.80 1.04 0.79
-Table 3: Upper limits at $95\%$ CL on the visible cross section $(\sigma\times A\times\epsilon)_{95}^{\rm obs}$ of the signal $pp\to\gamma+E_{T}^{\rm miss}$ for various signal regions (SRs) ATLAS:2020uiq .
+Table 3: Upper limits at $95\%$ CL on the visible cross section $(\sigma\times A\times\epsilon)_{95}^{\rm obs}$ of the signal $pp\to\gamma+E_{T}^{\rm miss}$ for various signal regions (SRs), where $A$ is the acceptance and $\epsilon$ is the efficiency ATLAS:2020uiq .
 Figure 9: Limit from ATLAS search ATLAS:2020uiq in the $(c_{3}^{\prime})_{UL}$ vs $M_{N_{2}}$ plane for signal region SRE1 (red) and SRI4 (black) for (a) $M_{\chi}=100$ GeV, $M_{N_{1}}=105$ GeV (b) $M_{\chi}=300$ GeV, $M_{N_{1}}=315$ GeV.
-In both cases $Y_{22}^{\prime}=10^{-9}\leavevmode\nobreak\ {\rm GeV}^{-1}$.
+In both cases $y_{22}^{\prime}=10^{-9}\penalty 10000\ {\rm GeV}^{-1}$.
 Figure 10: Upper limit on $c_{3}^{\prime}$ from the ATLAS search ATLAS:2020uiq in the $\delta_{1}-\delta_{2}$ plane for (a) $M_{\chi}=100$ GeV and (b) $M_{\chi}=300$ GeV.
-Here $Y_{22}^{\prime}=10^{-9}\leavevmode\nobreak\ {\rm GeV^{-1}}$.
+Here $y_{22}^{\prime}=10^{-9}\penalty 10000\ {\rm GeV^{-1}}$.
 The Lagrangian of this model is implemented in FeynRules(v2.3) Alloul:2013bka and the generated UFO files are used in the MC event generator MADGRAPH5(v3.4) Alwall:2014hca to simulate the signal events.
@@
 Fig.
-9 is for DM mass $M_{\chi}=100$ GeV, $M_{N_{1}}=105$ GeV and $Y_{22}^{\prime}=10^{-9}\leavevmode\nobreak\ \rm{GeV}^{-1}$.
+9 is for DM mass $M_{\chi}=100$ GeV, $M_{N_{1}}=105$ GeV and $y_{22}^{\prime}=10^{-9}\penalty 10000\ \rm{GeV}^{-1}$.
 Here the solid lines indicate current LHC limit while dashed lines are for HL-LHC projection.
@@
 Whereas, for SRE1, $\epsilon$ increases initially with $M_{N_{2}}$ and at a certain point it starts dropping.
-Therefor, $(c_{3}^{\prime})_{UL}$ varies differently for different SRs.
+Therefore, $(c_{3}^{\prime})_{UL}$ varies differently for different SRs.
 The limit is weakest when $M_{N_{2}}\simeq M_{N_{1}}$ as the photon becomes very soft.
@@
 Fig.
-9 shows the limits in the case $M_{\chi}=300$ GeV, $M_{N_{1}}=315$ GeV and $Y_{22}^{\prime}=10^{-9}\leavevmode\nobreak\ \rm{GeV}^{-1}$ .
+9 shows the limits in the case $M_{\chi}=300$ GeV, $M_{N_{1}}=315$ GeV and $y_{22}^{\prime}=10^{-9}\penalty 10000\ \rm{GeV}^{-1}$ .
 The dotted line represents $\Gamma(N_{2})/M_{N_{2}}=0.01$.
@@
 $M_{\chi}\,\,\rm{[GeV]}$  [$100$ , $500$]
-$\delta_{1}(M_{N_{1}}\leavevmode\nobreak\ \rm{[GeV]})$  $10^{-3}$ , $0.5$] $([\approx M_{\chi},750])$
-$\delta_{2}(M_{N_{2}}\leavevmode\nobreak\ \rm{[GeV]})$  $0^{-3} , 11$] $([\approx M_{N_{1}},1500])$
+$\delta_{1}(M_{N_{1}}\penalty 10000\ \rm{[GeV]})$  $10^{-3}$ , $0.5$] $([\approx M_{\chi},750])$
+$\delta_{2}(M_{N_{2}}\penalty 10000\ \rm{[GeV]})$  $0^{-3} , 11$] $([\approx M_{N_{1}},1500])$
 $y^{\prime}_{11}\,\,\rm{[GeV^{-1}]}$  $10^{-11}$ , $10^{-6}$]
@@
 4.
-Figure  13 shows that the mono-$\gamma$ search strongly constrain scenarios with large mass splitting between $M_{N_{1}}$ and $M_{N_{2}}$, these points are shown in red in the figure.
+Figure 13 shows that the mono-$\gamma$ search strongly constrain scenarios with large mass splitting between $M_{N_{1}}$ and $M_{N_{2}}$, these points are shown in red in the figure.
 Furthermore, as discussed above, the relic density constraint requires at least $N_{1}$ to be present in the early Universe, so that $\chi$ can be diluted through co-annihilation or inverse decay.
@@
 The dotted-color points in the $M_{N_{1}}$ and $M_{N_{2}}$ plane can be probed in HL-LHC while the red-crossed points remain out of HL-LHC reach.
-These points correspond to lower values of $M_{N_{2}}-M_{N_{1}}$ and/or lower value of $c^{\prime}_{3}$, which leads to suppressed signal efficiency and/or suppressed production cross section for the signal $pp\to N_{2}N_{2}\to\gamma+E_{T}^{\rm miss}$.
+These points correspond to lower values of $M_{N_{2}}-M_{N_{1}}$ and/or lower value of $c^{\prime}_{3}$, which leads to suppressed signal efficiency and/or suppressed production cross section for the signal $pp\to N_{1}N_{2}\to\gamma+E_{T}^{\rm miss}$.
 6 Summary and Conclusion
@@
 We emphasize the role of dimension-5 operators in the co-scattering mechanism.
-The key parameters on which DM dynamics depend are the masses $M_{N_{1,2}},M_{\chi}$ and couplings $\lambda,c_{3}^{\prime},Y_{11,22}$.
+The key parameters on which DM dynamics depend are the masses $M_{N_{1,2}},M_{\chi}$ and couplings $\lambda,c_{3}^{\prime},y^{\prime}_{11,22}$.
 Note that, we have assumed $c_{1,2}^{\prime}=0$ throughout our analysis.
@@
 As concerns the collider constraints, we find that the dimension-5 operators lead to sizable production of $N_{1,2}$ at the LHC and that the most stringent constraint is coming from the mono-$\gamma$ signature via $pp\to N_{1}N_{2}(\to N_{1}\gamma)\to\gamma+E_{T}^{\rm miss}$.
-Following the ATLAS analysis, we calculate the sensitivity of this signal which primarily depends on the mass difference $M_{N_{2}}-M_{N_{1}}=\delta_{2}$ (or the photon energy) and $c_{3}^{\prime}$.
+Following the ATLAS analysis, we calculate the sensitivity of this signal which primarily depends on the mass difference $M_{N_{2}}-M_{N_{1}}$ (or the photon energy) and $c_{3}^{\prime}$.
 We calculate the mono-$\gamma$ limit on $c_{3}^{\prime}$ in the $\delta_{2}-\delta_{1}$ plane, where $\delta_{1}$ dictates the relative mass splitting between $M_{N_{1}}$ and $M_{\chi}$.